OpenStudy (anonymous):

The sum of three integers is 47. If the first integer is doubled, the sum is 57. If the second integer is doubled, the sum is 62. Find the integers.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x+y+z=47 (1) 2x+y+z=57 (2) x+2y+z=62 (3) subtract (1) form (2) x=10 subtract (1) from (3) y=15 plug these into (1) 10+15+z=47 z=22 10,15,22

OpenStudy (anonymous):

Good job, but for simplicity, shouldn't you leave out z?

OpenStudy (anonymous):

you could do it by saying adding an extra x adds 10 so x=10, adding an extra y adds 15 so y is 15 then find the third. I was trying to turn the statements into equations literally I geuss.

OpenStudy (anonymous):

is that what you mean?